OpenStudy (anonymous):

Find the equation of f(x)=b^x given the point (2,e)

OpenStudy (id21):

Don't understand what exactly you want to find. You already have an equation, no? So only I can imagine is this x=2 and f(x)=e Then:\[e=b^2\] \[b=\pm \sqrt{e}=\pm1.65\]

OpenStudy (anonymous):

After getting b, I guess he just wants f(x) = sqrt(e)^x

OpenStudy (anonymous):

I need to get the equation in the format of f(x)=b^x given the point (2,e)

OpenStudy (anonymous):

That's it above in my answer.

OpenStudy (anonymous):

The point (2,e) allows you to deduce that b = e so f(x) = e^x (sorry, I put square root before).

OpenStudy (anonymous):

Blast, I had it right the first time. b is sqrt e so f(x) = sqrt(e)^x. Sorry.